40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 7 14:04:48 2011, steve, Update, SUS, He/Ne intensity noise effect on oplevs 7x
    Reply  Fri Oct 7 23:23:05 2011, kiwamu, Update, SUS, SRM oplev was oscillating 
       Reply  Wed Oct 19 10:49:35 2011, kiwamu, Update, SUS, SRM oscillation removed SRMoplevKWMtune.png
    Reply  Wed Oct 12 08:43:30 2011, steve, Update, SUS, ITMX oplev improved a bit ITMXoplev.pngITMXoplevservo_ON.png
       Reply  Thu Oct 13 14:23:09 2011, steve, Update, SUS, ITMX oplev with 3 mm beam on qpd ITMXoplev1103p#2.pngITMXoplev.png
          Reply  Mon Oct 17 17:07:30 2011, steve, Update, SUS, ETMX oplev returning beam od 3 mm ETMX-oplev.jpg
             Reply  Wed Oct 19 09:44:10 2011, steve, Update, SUS, ETMX oplev power spectrum ETMXoplev2lens.png
             Reply  Fri May 18 15:25:15 2012, steve, Update, SUS, AR coated lens swapped in at ETMX oplev ETMXoplevlensAR.png
Message ID: 5660     Entry time: Thu Oct 13 14:23:09 2011     In reply to: 5655     Reply to this: 5680
Author: steve 
Type: Update 
Category: SUS 
Subject: ITMX oplev with 3 mm beam on qpd 

 I replaced the JDSU-Uniphase 1125P by 1103P He/Ne laser. This new laser had 2.8 mW output yesterday. It degraded to 0.5 mW by this morning.

The beam size on the QPD is ~3 mm  This should give us  better sensitivity. These are not the perfect lenses at all, but we have them here.

On the other hand, there are still some coherence below 1 Hz, so the laser intensity noise or clipping dominating  this  part of the spectrum.

 

Attachment 1: ITMXoplev1103p#2.png  51 kB  | Show | Hide all | Show all
Attachment 2: ITMXoplev.png  2.731 MB  Uploaded Thu Oct 13 15:25:00 2011  | Hide | Hide all | Show all
ITMXoplev.png
ELOG V3.1.3-